OpenStudy (anonymous):

Evaluate: 3.5-2x for x=2.4

OpenStudy (anonymous):

3.5 - 2*(2.4) = ?

OpenStudy (anonymous):

1.3

OpenStudy (anonymous):

... 3.5 is less than 4.8 so its negative.

OpenStudy (mathstudent55):

3.5 - 2*(2.4) = 3.5 - 4.8 4.8 is greater than 3.5, so when you subtract answer must be negative.

OpenStudy (anonymous):

so its -1.3

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks